likeabossssssss (likeabossssssss):

he sum of any two consecutive integers is a composite number plzhelp

likeabossssssss (likeabossssssss):

a 16+17=33 b 10+11=21 c 6+7+13 d 7+8=15

OpenStudy (danjs):

composite number

OpenStudy (danjs):

omposite number is a positive integer that has at least one positive divisor other than one or the number itself. Any integer greater than one that is not a prime number is a composite number,[1][2] while the number one is a unit; it is neither prime nor composite from wikipedia

likeabossssssss (likeabossssssss):

ok

likeabossssssss (likeabossssssss):

idont think i spelled it right

OpenStudy (danjs):

sum of 2 consecutive integers -- n + (n + 1) for n is an integer

likeabossssssss (likeabossssssss):

i think it is c wbu

OpenStudy (danjs):

all 4 choices have sums of consecutive integers on the left

OpenStudy (danjs):

only c) 13 is a prime number, and not a composite number

OpenStudy (danjs):

a,b, and d are all correct

likeabossssssss (likeabossssssss):

i think it is c wbu

OpenStudy (danjs):

21 = 7*3 33 = 3*11 15 = 5*3 for example, all those are composite does it ask which one is not true? then c all others are good

likeabossssssss (likeabossssssss):

it say it has to end in a prime number